a cabinet, 4 chairs and 2 tables cost $825. The cabinet costs 3 times as much as a chair. The table costs $120 more than a chair. What is the cost of the cabinet.
1
Expert's answer
2014-06-12T03:39:53-0400
Assume that a chair costs x$. Then the cabinet costs 3x$.As table costs $120 more than a chair, then it costs x+120$. We know that a cabinet, 4 chairs and 2 tables cost 825$. We can write that as 3x+4*x+2(x+120)=825. We obtain 9x=585$, or x=585/9$
So, the cabinet costs 3x=3*585/9=585/3$=195$.


Answer. 195$